# Predicting Maintenance Cessation of Open Source Software Repositories with An Integrated Feature Framework

### v1

@@ Abstract.
The maintenance risks of open source software (OSS) projects pose significant threats to the quality, security, and resilience of modern software supply chains. While prior research has proposed diverse approaches for predicting OSS maintenance risk—leveraging signals ranging from surface features (e.g., stars, commits) to social network analyses and behavioral patterns—existing methods often suffer from ambiguous operational definitions, limited interpretability, and datasets of insufficient scale or generalizability. In this work, we introduce “maintenance cessation”, grounded in both explicit archival status and rigorous semantic analysis of project documentation. Building on this foundation, we curate a large-scale, longitudinal dataset of 115,466 GitHub repositories—encompassing 57,733 confirmed cessation events—complemented by comprehensive, timeline-based behavioral features. We propose an integrated, multi-perspective feature framework for predicting maintenance cessation, systematically combining user-centric features, maintainer-centric features and project evolution features. AFT survival analysis demonstrates a high C-index (0.846), substantially outperforming models relying only on surface features. Feature ablation and SHAP analysis further confirm the effectiveness and interpretability of our approach. Finally, we demonstrate real-world applicability by deploying a GBSA classifier in the openEuler ecosystem for proactive package risk screening. Our work establishes a scalable, interpretable foundation for maintenance-risk prediction, enabling reproducible risk management across large-scale open source ecosystems.
@@ 1. Introduction
Meanwhile, the growing dependence on open source software means the growing risk of ecosystem-wide disruptions. The halt of maintenance is not rare in the open source ecosystem, and it poses the sudden risk of software defects, security vulnerabilities, and more to numerous downstream users and dependent software systems (Kula et al., 2017). The OSSRA report finds that 91% of audited OSS components exhibited no clear signs of maintenance in the past two years (Synopsys, 2025), highlighting the prevalence of this issue. And notable, high-profile OSS projects may also run into discontinuation—well-known examples include Atom (Atom, [n. d.]), Brackets (adobe, [n. d.]), and faker.js (Marak, [n. d.]).
Proactive prediction of unmaintainance in OSS projects is a promising approach to mitigate this risk. By warning developers and organizations about the unmaintenance risk, it enables them to take proactive countermeasures including vendoring, removing, or migrating risky dependencies to anticipate disruptions. While extensive research has explored the potential of data-driven unmaintenance risk prediction (Samoladas et al., 2010a; Xia et al., 2022), two gaps must be addressed for the prediction techniques to be practical.
The first gap lies in the definition and identification of unmaintenance. Many of previous studies of project “deprecation” or “failure” defined unmaintenance based on the activity features, they often assume a project is abandoned after a fixed threshold of inactivity (Coelho and Valente, 2017; Liao et al., 2023; Robinson et al., 2022; Malviya Thakur et al., 2025). However, we found counter-examples of “revival” in real-life open source projects to their assumption (Section 3.1), indicating that inactivity is not a reliable proxy of project unmaintainance.
@@
The second gap is in the prediction techniques. Existing approaches for prediction typically use surface features such as the number of star, commit, or pull request (PR), which are volatile, susceptible to manipulation (He et al., 2024), and insufficient for capturing deeper sustainability dynamics such as governance transitions, contributor attrition, or changes in user engagement. Other approaches that incorporate more high-level features and advanced methods often lack semantic interpretability, which questions their practical reliability and generalizability.
In this paper, we introduce the term maintenance cessation as a clear and definitive definition of the end of project maintenance. Specifically, we identify maintenance cessation based on the union of two criteria: (1) explicit GitHub archived status (GitHub Documentation, 2025; GitHub, nd), and (2) unambiguous maintenance cessation statements in project documentation (Coelho and Valente, 2017). Cessation is both semantically precise and task-specific, and maintenance cessation more closely captures the genuine discontinuation of maintenance, explicitly excludes projects that are merely “quiet” or low-activity but remain responsibly maintained (Naberezny et al., 2011; Ashkenas et al., 2013; Suuronen et al., 2022).
@@
Based on the definition of maintenance cessation, we curate a large-scale longitudinal dataset of 115,466 GitHub repositories, including 57,733 with validated maintenance cessation events and comprehensive behavioral timelines (Section 3.3). We then propose a multi-dimensional, interpretable feature framework that goes beyond surface features by jointly modeling: (1) user influence using a time-aware, weighted bipartite user-project interaction network; (2) maintainer behavior, including activity recency and response latency; and (3) project evolution characteristics, such as community role balance and the shift between feature and bugfix contributions (Section 4.1). Using an accelerated failure time (AFT) survival model, we empirically demonstrate that our approach substantially outperforms existing methods, while providing improved interpretability and early warning. We further train a GBSA classifier and integrated it into openEuler’s risk assessment pipeline (Section 5). It proves that our framework is practical in assisting the risk control and management of real-life large-scale software supply chains.
To sum up, in this paper, we:
  • Proposed a precise definition of OSS maintenance cessation.
  • Constructed a large-scale, longitudinally labeled dataset of OSS repositories.
  • Designed a multi-dimensional, interpretable feature framework for predicting maintenance cessation.
  • Empirically demonstrated substantial predictive improvements over existing methods.
  • Validated the real-world utility of our approach through deployment in the openEuler ecosystem.
We believe our approach lays a crucial foundation for the future of software supply chain risk assessment, OSS dependency management, and sustainable ecosystem governance.
@@ 2. Background
The challenge of predicting when an open source project will cease to be maintained—whether referenced as “abandonment,” “deprecation”,“failure”, or more broadly as sustainability—has generated extensive research attention in the software engineering domain. Early work in this area primarily adopted statistical survival analysis approaches drawn from reliability engineering to model project longevity and risk. For example, models such as the Cox proportional hazards model and the Kaplan-Meier estimator were used to investigate how factors like the number of active contributors and the presence of major version releases correlated with continued OSS project maintenance or survival (Samoladas et al., 2010a; Robinson et al., 2022; Khondhu et al., 2013; Samoladas et al., 2010b; Li et al., 2022a; Valiev et al., 2018; Mujahid et al., 2022).
With the increasing availability of large-scale data from social coding platforms (e.g., GitHub), more recent research has turned to machine learning (ML)-based methods. Most existing approaches operationalize “abandonment” or “deprecation” based on symptoms such as prolonged inactivity in commits, issues, or PR, and less often, declarations from maintainers in documentation (Liao et al., 2023; Moid et al., 2021; Li et al., 2022b). Feature engineering in these works has focused primarily on surface-level, readily available project activity statistics—counts of stars, commits, issues, PRs, and sometimes doc structure or languages. Some research has explored augmenting these signals with social or network-derived features, such as developer influence or collaboration centrality measures, or has employed more advanced architectures including ensemble boosting and graph neural networks (Ren et al., 2020; Liao et al., 2023; Xia et al., 2022; Coelho et al., 2018, 2020).
Despite these advances, several key limitations hamper the robustness, interpretability, and practical adoption of the current state-of-the-art. Specifically, we identify several major areas where existing research falls short:
Ambiguous or oversimplified definitions of maintenance cessation. The majority of prior work equates “abandonment” or “deprecation” with simple inactivity over a fixed period—typically time since last commit or last issue. However, these terms are ambiguous in meaning, and ignores the reality that many mature projects may operate with infrequent updates but still provide essential maintenance and user support, while some genuinely unmaintained projects might not be captured by activity alone. The lack of a universally precise and robust definition for true maintenance cessation.
@@
Recent advances have begun to address some of these shortcomings, but major gaps persist, particularly in integrating behavioral signals, scaling to large real-world ecosystems, and providing definitions and methods that are both theoretically sound and operationally actionable.
3. Methodology
This section presents our unified methodological framework for predicting maintenance cessation of OSS repositories. Our approach encompasses formal problem definition, a scalable and rigorous pipeline for data construction and labeling, the design of interpretable and discriminative feature sets characterizing user, maintainer, and project dynamics, and robust predictive modeling strategies tailored to both research and operational deployment.
3.1. Problem Definition
Accurately predicting when an OSS project will cease to be actively maintained is a critical but challenging task for both researchers and practitioners. However, before prediction is possible, a precise and actionable definition of “maintenance cessation” is essential. Existing literature uses various terms such as “abandonment”, “deprecation” or “failure”, but often with vague meaning and inconsistent operationalization: some studies rely solely on indicators of prolonged inactivity (e.g., the absence of commits within a fixed period) (Coelho and Valente, 2017; Liao et al., 2023; Robinson et al., 2022; Malviya Thakur et al., 2025), while others require declarations from maintainers in documentation or metadata. Moreover, these terms carry domain-specific meanings in other areas of software engineering, which can lead to semantic confusion (Robbes et al., 2012; Haryono et al., 2021; Song et al., 2024). This ambiguity causes significant challenges for scientific reproducibility and practical application:
Over-reliance on (in)activity: Many mature OSS projects exhibit infrequent updates but remain under responsible maintenance. For example, supervisor (Naberezny et al., 2011), backbone (Ashkenas et al., 2013), and underscore.string (Suuronen et al., 2022) may go several months—or even over a year—without any commits. However, their maintainers still assume stewardship over the repositories, occasionally performing code modifications when critical issues arise. Treating such low-frequency activity as a sign of unmaintained can result in high false-positive rates.
Semantic confusion: Terms such as “abandonment,” “deprecation,” or “failure” may refer to ceasing feature development, stopping issue triage, discontinuing support for older versions, or ending all development and support. Without explicit definition and consistent definition, label quality and downstream prediction suffer.
In this work, we adopt the term maintenance cessation to describe the precise condition in which the maintainers have explicitly and definitively declared the end of maintenance for a software repository. This term is deliberately chosen not merely to avoid the ambiguity of commonly used alternatives like abandonment or deprecation, but because it clearly and unambiguously communicates the specific event we aim to detect-namely, an intentional, declared termination of maintenance. Compared to other vague or overloaded terms, maintenance cessation is immediately interpretable and directly aligned with our task objective. It emphasizes the cessation of active stewardship, without presuming disuse, obsolescence, or removal from the ecosystem. Our definition explicitly excludes projects that are simply low-activity, stable, or “quiet” but are still under responsible maintenance.
Accordingly, we identify maintenance cessation as occurring if, and only if, at least one of the following two conditions is satisfied:
@@ 3.1. Problem Definition
    Explicit archival status: The repository is marked as “archived” on GitHub, which irreversibly freezes the repository in a read-only mode and reflects a maintainer’s explicit cessation of maintenance (GitHub, nd; GitHub Documentation, 2025). The timestamp of archiving is recorded as the cessation event.
@@
    Semantic cessation statement: The repository’s top-level documentation (e.g., README, project description) contains direct, unambiguous textual statements indicating that the repository is deprecated, no longer maintained, unmaintained, or has reached end-of-life. To minimize ambiguity, we detect such statements through a hybrid pipeline that combines targeted keyword filtering, manual verification of a large sample, and large language model (LLM) and transformer-based classification, as detailed in Section 3.3.
This dual-criteria definition ensures: (1) precision—capturing only verifiable maintenance cessation events, not mere inactivity, and (2) recall—identifying cases explicitly signaled in text but not only using GitHub’s archival feature. These clear labels form the basis for predictive modeling in the remainder of this work.
3.2. Methodology Overview
Figure 1 provides an overview of our methodological framework for predicting maintenance cessation of OSS repositories. Our approach is structured around five core components, each tightly integrated to ensure scientific rigor, feature interpretability, and real-world applicability.
Figure 1. Methodology Overview.
@@ 3.2. Methodology Overview
    Defining Maintenance Cessation. We establish a dual-criteria definition where repositories cease maintenance if explicitly archived or containing semantic cessation declarations.
@@
    Data Construction. We collect GitHub repositories (stars $\geq$32) and assign cessation labels through hybrid annotation combining keyword filtering, manual validation, and automated classification.
@@
    Feature Engineering. We develop a multi-perspective framework including surface features, user influence scores from bipartite networks, maintainer behavioral signals, and project evolution indicators.
  4. (4)
    Feature Effectiveness Validation. We validate features through survival analysis using AFT models with ablation studies and SHAP interpretability, supplemented by temporal case analyses.
  5. (5)
    Real-World Deployment. We implement risk prediction in openEuler ecosystem via GBSA classifiers applied to upstream package mappings.
By integrating precise definitions, scaleable data processing, interpretable feature frameworks, rigorous modeling, and operational validation—as visually summarized in Figure 1—our methodology delivers a reliable, reproducible, and actionable blueprint for large-scale OSS maintenance cessation prediction.
3.3. Data Construction and Preprocessing
To support robust, generalizable prediction and facilitate downstream research, we constructed a comprehensive, high-quality dataset capturing both longitudinal behavioral traces and rigorously validated cessation labels for a large population of OSS repositories. This process comprised repository selection, multi-source data collection, careful labeling, temporal aggregation, and quality control. The Data Construction and Preprocessing box in Figure 1 shows the general process.
Repository Selection and Scope.
We target repositories hosted on GitHub (GitHub, 2007), the world’s largest collaborative development platform. To balance representativeness and tractability, we include all non-fork repositories with at least 32 stars, following both prior studies and API limit considerations. This threshold effectively filters out trivial or inactive repositories, while preserving coverage across diverse languages, domains, and activity levels. Our time span covers repositories created from 2011 through 2025, ensuring sufficient history for meaningful temporal analysis. Forked projects are excluded to avoid redundancy and ambiguous maintenance signals linked to upstream dependencies.
Ground Truth Labeling: Maintenance Cessation.
Accurately labeling the cessation of software maintenance poses a significant challenge due to its informal and distributed nature. We adopt a three-step pipeline to construct high-quality, reproducible ground truth labels for repository maintenance cessation:
  1. (1)
    Archived Repository Collection. We leverage the GitHub GraphQL API (Github, [n. d.]) to retrieve repositories explicitly marked as ”archived”. This status, once enabled by maintainers, freezes the repository in a read-only mode and is treated as a strong, platform-verified signal of maintenance cessation. To reduce noise from trivial or inactive projects, only repositories with more than 32 stars are retained, aligning with both practical and API efficiency considerations (Github, [n. d.]). The archival timestamp is recorded as the official cessation date.
  2. (2)
    Cessation Declaration Mining. As prior work (Coelho and Valente, 2017) has shown, many developers choose to declare “deprecation” or “abandonment” in README files or repository descriptions rather than use the archive flag. We identify such cases using a semantic statement extraction strategy: (1) A list of maintenance cessation keywords (e.g., “no longer maintained”, “deprecated”, “unmaintained”, etc.) is first used to retrieve candidate repositories. To mitigate false positives due to ambiguous usage (e.g., deprecating a dependency rather than the project itself), we perform manual annotation on 1,200 stratified samples. (2) The annotation task is conducted using Label Studio (HumanSignal, 2024), where two independent annotators label each sample based on both textual signals and repository activity metadata (e.g., commits, issues, PRs). The resulting labels achieve a Cohen’s kappa of 0.814, indicating substantial inter-annotator agreement. (3) Using 600 samples for training and 600 for validation, we fine-tune a SetFit classifier (Tunstall et al., 2022) based on sentence transformers (Reimers and Gurevych, 2019). Our best model, built on the paraphrase-mpnet-base-v2 checkpoint (sentence-transformers, 2022), achieves an accuracy of 0.96, recall of 0.96, and precision of 0.90.
  3. (3)
    LLM-Assisted Labeling. To scale high-precision labeling, we evaluated multiple large language models (LLMs) on the 1,200-sample benchmark. Among the tested models (including GPT-4o, DeepSeek variants, Claude 3.7, and Qwen series), GPT-4o demonstrated superior performance with accuracy, precision, recall, and F1 achieving 0.9463, 0.9429, 0.8985, and 0.9202. This model was consequently selected to label the remaining 37,065 samples. We observed 95.91% agreement between GPT-4o and our SetFit classifier. All repositories where GPT-4o and SetFit predictions disagreed were manually reviewed and annotated by authors to resolve discrepancies and ensure final label quality.
This process results in 57,733 repositories with confidently labeled maintenance cessation dates, derived from either archival flags or timestamped textual cessation declarations.
@@ 3.3. Data Construction and Preprocessing
To capture project evolution and enable feature engineering, we aggregate fine-grained, time-series records of repository activities using a combination of GHArchive event data (GHArchive, 2024) and the GitHub API. For each repository, we record (per month):
@@
Compared with other public datasets (e.g., GHTorrent (Gousios, 2013)), GHArchive is leveraged for its completeness and efficiency in reconstructing historical events at scale. Events are parsed using distributed ETL pipelines and loaded into a column-oriented OLAP system (ClickHouse), supporting performant querying and flexible aggregation.
Figure 2. Maintenance lifespan distribution of ceased repositories.
Data Cleaning and Quality Control.
Quality control steps include: (1) Removal of repositories with inconsistent, corrupted, or incomplete event histories. (2) Cross-validation of label assignments to mitigate false positives, e.g., distinguishing actual project maintenance cessation from critical version maintenance cessation or temporary inactivity. (3) Statistical analysis of activity, star, and lifetime distributions for ongoing vs. ceased repositories, confirming representativeness and detecting potential annotation or sampling bias (see Table 1 and Figure 2).

### v2

# Predicting Abandonment of Open Source Software Projects with An Integrated Feature Framework

@@ Abstract.
Open Source Software (OSS) is a cornerstone of contemporary software development, yet the increasing prevalence of OSS project abandonment threatens the quality, security, and resilience of global software supply chains. Although previous research has explored a range of abandonment prediction methods—drawing on signals from simple activity metrics (e.g., stars, commits) to social network and behavioral analyses—these methods often demonstrate unsatisfactory predictive performance, further plagued by imprecise abandonment discrimination, limited interpretability, and a lack of large, generalizable datasets. In this work, we address these challenges by reliably detecting OSS project abandonment through a dual approach: explicit archival status and rigorous semantic analysis of project documentation or description. Leveraging a precise and scalable labeling pipeline, we curate a comprehensive longitudinal dataset of 115,466 GitHub repositories, encompassing 57,733 confirmed abandonment repositories, enriched with detailed, timeline-based behavioral features. Building on this foundation, we introduce an integrated, multi-perspective feature framework for abandonment prediction, capturing user-centric, maintainer-centric, and project evolution features. Survival analysis using an AFT model yields a high C-index of 0.846, substantially outperforming models confined to surface features. Further, feature ablation and SHAP analyses confirm both the predictive power and interpretability of our approach. We further demonstrate practical deployment of a GBSA classifier for predicting package risk in the openEuler ecosystem. By unifying precise labeling, multi-perspective features, and interpretable modeling, our work provides reproducible, scalable, and practitioner-oriented support for understanding and managing abandonment risk in large OSS ecosystems. Our tool not only predicts abandonment but also enhances program comprehension by providing actionable insights into the health and sustainability of OSS projects.
@@ 1. Introduction
Meanwhile, the growing dependence on OSS means the growing risk of ecosystem-wide disruptions. OSS projects are typically maintained by a small number of developers on a voluntary basis. These developers may discontinue their involvement in project maintenance for various reasons, such as the project no longer meeting public needs, being replaced by competing alternatives, loss of personal interest, and so forth (Coelho and Valente, 2017). Once no one continues to maintain the project, the OSS is considered abandoned, which is not rare in the open source ecosystem, and it poses sudden risks of software defects, security vulnerabilities, and more to numerous downstream users and dependent software systems (Kula et al., 2017). The OSSRA report finds that 91% of audited OSS components exhibited no clear signs of maintenance in the past two years (Synopsys, 2025), highlighting the prevalence of this issue. Notably, high-profile OSS projects may also run into abandonment—well-known examples include Atom (Atom, [n. d.]), Brackets (adobe, [n. d.]), and faker.js (Marak, [n. d.]).
Predicting OSS project abandonment is crucial not only for risk mitigation, but also for supporting practical program comprehension tasks, such as dependency analysis and architectural assessment. Interpretable risk signals allow developers to better understand and manage the sustainability and evolution of their software dependencies, leading to more robust and informed decision-making. While extensive research has explored the potential of data-driven abandonment prediction (Samoladas et al., 2010a; Xia et al., 2022; Liao et al., 2023), two challenges must be addressed for these prediction techniques to be practical.
The first challenge lies in the discrimination of abandonment. Some previous studies have defined project abandonment solely based on activity metrics. They often assume a project is abandoned after a fixed threshold of inactivity (Robinson et al., 2022; Malviya Thakur et al., 2025). We found counter-examples of “revival” in real-life non-trivial OSS projects to their assumption (Naberezny et al., 2011; Ashkenas et al., 2013; Suuronen et al., 2022), indicating that inactivity cannot definitively determine abandonment. Several prior works have employed explicit abandonment statements as the criterion to discriminate abandonment, and our method also adopts this approach for determination. Notably, previous studies (Liao et al., 2023) often used direct keyword matching in project descriptions for discrimination. Within our manually labeled set of 1,174 randomly selected keyword-matched OSS repositories, only 404 (34.4%) had README files or descriptions that genuinely indicated the repository was abandoned. The remaining 768 repositories (65.6%) contained no explicit statement of abandonment. Both SetFit and LLM methods, and prior research (Coelho and Valente, 2017) yielded similar results, indicating that the previously used keyword-matching method for identifying abandonment contained a substantial proportion (over 60%) of false positives—a figure that cannot be overlooked. These repositories, lacking genuine abandonment statements, likely triggered the keyword matches for various reasons, such as declaring a specific version abandoned, stating that another project was abandoned, or indicating that a particular component was abandoned. Some studies (Coelho and Valente, 2017; Miller et al., 2023, 2025) have used manual labelling for accurate identification of abandonment statements in text, albeit often on a small scale due to the intensive labor involved.
@@
The second challenge concerns predictive features. Existing approaches typically use surface features such as the number of star, commit, or pull request (PR), which are volatile, susceptible to manipulation (He et al., 2024), and insufficient for capturing deeper sustainability dynamics such as governance transitions, contributor attrition, or changes in user engagement. Other approaches that incorporate more high-level features and advanced methods often lack interpretability, such as straightforward feeding information into GNNs or other complex models (Liao et al., 2023). These features fail to support genuine program comprehension regarding a project’s health and sustainability. Often, they merely produce a numeric value whose semantic meaning is difficult to interpret and actionable insights for maintainers remain unclear (Faber et al., 2021). Regarding some of the more interpretable advanced features proposed in existing studies, certain limitations remain; we provide a detailed discussion in Section 4.
Project abandonment is identified through a union of two criteria: (1) explicit GitHub archived status (GitHub Documentation, 2025; GitHub, nd), and (2) unambiguous abandonment statements in project documentation or description (Coelho and Valente, 2017). We implemented this through a precise pipeline that combines manual labeling, ML, and LLM. This approach virtually eliminates the false positives, which accounted for over 60% of labels in prior keyword-matching methods, while enabling rapid processing of large data volumes to construct a large-scale dataset.
@@
From this precise labelling, we have generated a large-scale longitudinal dataset of 115,466 GitHub repositories, including 57,733 with validated abandonment events and comprehensive behavioral timelines (Section 3.2). We then propose a multi-dimensional, interpretable feature framework that goes beyond surface features by jointly modeling: (1) user influence via a time-aware, weighted bipartite user-project network; (2) maintainer behavior, including activity recency and response latency; and (3) project evolution characteristics, such as community role balance and feature/bugfix contribution shifts (Section 4.1). Using an accelerated failure time (AFT) model, our approach achieves a high C-index (0.846), filling a critical gap in existing research while providing superior predictive power, interpretability, and early warning capabilities. Beyond predictive accuracy, our feature framework enables developers and ecosystem stakeholders to better comprehend the underlying causes, behavioral signals, and community dynamics of OSS project abandonment. Furthermore, to demonstrate real-world applicability, we deployed a GBSA classifier trained on our framework within the openEuler ecosystem to conduct proactive risk screening of its packages (Section 5). This demonstrates the framework’s practical utility in assessing risks for large-scale, real-world software supply chains.
To sum up, in this paper, we: (1) Proposed a precise method for discrimination true abandonment; (2) Constructed a large-scale, longitudinally labeled dataset of OSS repositories; (3) Designed a multi-dimensional, interpretable feature framework for predicting abandonment; (4) Empirically demonstrated superior predictive performance; (5) Validated the real-world utility of our approach through deployment in the openEuler ecosystem.
By establishing a rich foundation for program comprehension and risk awareness, our work contributes to software supply chain risk assessment, OSS dependency management, and sustainable ecosystem governance. As illustrated in Figure 1, this foundation is built upon a methodological framework comprising four core components, which are tightly integrated to ensure scientific rigor, feature interpretability, and real-world applicability.
Figure 1. Methodology Overview.
@@ 2. Background
Predicting open source project abandonment has attracted significant research interest in software engineering. Early work primarily adopted survival analysis methods from reliability engineering, using models like Cox proportional hazards and Kaplan-Meier estimators to examine how factors such as active contributor count and major version releases correlate with project survival (Samoladas et al., 2010a; Robinson et al., 2022; Khondhu et al., 2013; Samoladas et al., 2010b; Li et al., 2022a; Valiev et al., 2018; Mujahid et al., 2022).
With the growth of social coding platforms, recent studies have shifted toward machine learning methods. Most define abandonment through symptoms like prolonged inactivity in commits or pull requests, or occasionally through maintainer declarations in documentation (Liao et al., 2023; Moid et al., 2021; Li et al., 2022b; Miller et al., 2023, 2025). Feature engineering in these works relies heavily on surface-level activity counts—stars, commits, issues, PRs—sometimes supplemented with social or network features, or advanced architectures such as GNNs and ensemble methods (Ren et al., 2020; Liao et al., 2023; Xia et al., 2022; Coelho et al., 2018, 2020).
Despite these advances, several key limitations hamper the robustness, interpretability, and practical adoption of the current state-of-the-art:
Imprecise Abandonment Criteria. The absence of precise abandonment criteria in most previous work leads to high false-positive rates. For example, keyword-matching mislabels over 60% of repositories, and definitive signals like GitHub’s ”archived” status are often overlooked.
@@
While recent efforts address some gaps, challenges remain in scaling interpretable methods, improving abandonment discrimination, and enabling actionable risk assessment in production environments.
3. Foundations for Prediction
This section details the foundational methodology for our study on OSS project abandonment prediction. We first address the critical prerequisite of operationalizing abandonment through a precise, dual-criteria definition. We then describe the construction of the large-scale, longitudinal dataset that underpins all subsequent analysis and modeling in this work.
3.1. Operationalization of Abandonment
Accurately predicting OSS project abandonment is critical yet challenging. A prerequisite for accurate prediction is the precise identification of abandonment events. Existing research shows inconsistent operationalizations: some rely solely on prolonged inactivity (e.g., no commits within a fixed period) (Liao et al., 2023; Robinson et al., 2022; Malviya Thakur et al., 2025), while others use keyword matching to detect maintainer declarations (Coelho and Valente, 2017; Liao et al., 2023). Unfortunately, keyword-based methods incur high false-positive rates, as confirmed in our study, while precise manual labeling does not scale to large datasets.
In this work, we define a project as abandoned if at least one of the following conditions is met:
@@ 3.1. Operationalization of Abandonment
    Explicit archival status: The repository is marked as “archived” on GitHub, making it read-only and signaling explicit maintainer abandonment (GitHub, nd; GitHub Documentation, 2025). The archiving timestamp defines the abandonment time.
@@
    Semantic statement of abandonment: The project’s top-level documentation contains clear textual statements (e.g., “abandoned,” “no longer maintained,” “deprecated,” or “end-of-life”). We detect these using a hybrid pipeline combining keyword filtering, manual verification, machine learning and LLM-based classification, as detailed in Section 3.2.
This dual-criterion approach ensures both recall (capturing text-based signals beyond archiving) and precision (eliminating over 60% of false positives from keyword matching alone). The resulting high-quality labels underpin all subsequent predictive modeling in this work.
3.2. Data Construction and Preprocessing
To support robust and generalizable prediction, we constructed a comprehensive dataset capturing longitudinal behavioral traces and rigorously validated abandonment labels for a large population of OSS repositories. The construction process included repository selection, multi-source data collection, careful labeling, temporal aggregation, and quality control, as outlined in the Data Construction and Preprocessing box of Figure 1.
Repository Selection and Scope.
We focus on non-fork repositories hosted on GitHub (GitHub, 2007), the world’s largest collaborative development platform. To balance representativeness and tractability, we include all repositories created between 2011 and 2025 with at least 32 stars, following prior studies (Xiao et al., 2023) and API limit (Github, [n. d.]) considerations. This threshold filters out trivial or inactive repositories, while preserving diversity in languages, domains, and activity levels.
Ground Truth Labeling.
We implement a three-step pipeline to produce high-quality, reproducible abandonment labels:
@@ 3.2. Data Construction and Preprocessing
    Archived Repository Collection. Using the GitHub
    GraphQL API (Github, [n. d.]), we collect repositories explicitly marked as ”archived” — a read-only state treated as a strong, platform-verified abandonment signal. The archival timestamp is recorded as the abandonment date.
@@
    Abandonment Declaration Mining. Many developers choose to declare abandonment in README files or repository descriptions rather than use the archive flag (Coelho and Valente, 2017; Miller et al., 2023, 2025). We identified such cases using a semantic statement extraction strategy: (1) Following prior works (Liao et al., 2023; Coelho and Valente, 2017; Miller et al., 2025), we first use a list of abandonment keywords (e.g., “abandoned”, “deprecated”, “unmaintained”, etc.) is first used to retrieve candidate repositories. To mitigate false positives due to ambiguous usage (e.g., abandon a dependency rather than the project itself), we perform manual annotation on 1,174 (3%) randomly selected samples. (2) The annotation task is conducted using Label Studio (HumanSignal, 2024), where two authors label each sample based on textual signals. The first round of labeling ended with an agreement of 95% and a Cohen’s kappa of 0.814, which suggests a strong agreement. All inconsistencies were resolved through discussion, leading to a final consensus on the entire dataset. (3) Using 587 samples for training and 587 for validation, we fine-tune a SetFit classifier (Tunstall et al., 2022) based on sentence transformers (Reimers and Gurevych, 2019). Our best model, built on the paraphrase-mpnet-base-v2 checkpoint (sentence-transformers, 2022), achieved 0.96 accuracy, 0.96 recall, and 0.90 precision.
@@
    LLM-Assisted Labeling. To scale high-precision labeling, we evaluated multiple large language models (LLMs) on the 3%-sample manual labelling benchmark. Among the tested models (including GPT-4o, DeepSeek-V3, DeepSeek-R1, Claude-3.7, and Qwen series), GPT-4o performed best with accuracy, precision, recall, and F1 achieving 0.9463, 0.9429, 0.8985, and 0.9202. This model and SetFit was consequently selected to label remaining samples. We observed 95.91% agreement between GPT-4o and our SetFit classifier. All disagreements were resolved through discussion, resulting in 1,515 relabeled samples during manual cross-validation.
This process yields 57,733 repositories with confidently labeled abandonment dates from archival flags or textual declarations, eliminating all the 60%+ false positive rate observed in prior keyword-based methods.
@@
To capture project evolution, we aggregate time-series repository activities using
GHArchive (GHArchive, 2024) and GitHub API. Monthly records include:
@@
Compared with other public datasets (e.g., GHTorrent (Gousios, 2013)),
GHArchive is leveraged for its completeness and efficiency in reconstructing historical events at scale. Events are parsed via distributed ETL and stored in ClickHouse for efficient querying.
Quality Control and Initial Analysis.
Our quality control process began with the removal of repositories exhibiting inconsistent, corrupted, or incomplete event histories, typically resulting from missing API data or API reconciliation failures. This initial filtering step affected a negligible portion (0.3%) of the dataset, ensuring data integrity without materially influencing the overall results.
Initial statistical analysis compared star counts and lifespans between active and abandoned repositories to validate dataset representativeness and screen for potential annotation biases. Figure 2 shows the lifespan distribution of 23,730 abandoned projects, averaging nearly four years with high variability—underscoring the difficulty of predicting abandonment.
Figure 2. Lifespan distribution of abandoned repositories.
Further insights are drawn from the star count distributions in Table 1. Abandoned projects demonstrate significantly lower mean and maximum star counts compared to the full population, consistent with the expectation that abandoned projects attract diminishing attention. The considerably smaller standard deviation and maximum star counts further indicates that abandoned projects lack the highly popular repositories found in the overall dataset.
Notably, abandoned projects show a higher median star count. Abandonment is discriminated through explicit maintainer actions, which are more likely to occur in projects that have already garnered a moderate level of community attention. Consequently, repositories abandoned silently with minimal user engagement are under-represented in our labeled abandoned cohort, slightly inflating the median star count for confirmed abandoned projects. This initial analysis validates our data quality and proves that our data is representative.